On January 1, 2019, the Leaf Company purchased a patent for 70 000
The patent has a legal life of 20 years and an estimated useful life of 10 years
INSTRUCTIONS: Prepare the journal entry to record 2019 amortization expense.
2)
On October 1, 2019, Beaufort Mining Inc. purchased a new mine.
Cost of the mine 640 000
Estimated tons of ore in this mine 500 000 tons
Ore mined and sold during 2017 80 000 tons
INSTRUCTIONS: Prepare the journal entry to record depletion expense for 2019
Answer #1

1. Patent:

A patent is amortized over the shorter of its estimated useful life or legal life. Here, legal life is 20 years and estimated useful life is 10 years. So, the patent will be amortized over 10 years (being shorter than legal life).

Amortization expense for 2019 = Cost / Useful life

Amortization expense = $70000 / 10 = $7000

Required journal entry for amortization expense for 2019 is:

Debit Amortization expense $7000

Credit Accumulated Amortization    $7000

2. Depletion expense:

Under depletion, first we will calculate the unit depletion rate and then we will calculate the depletion expense.

Unit depletion rate = Depletion base / Estimated recoverable units

where, Depletion base = Cost - Salvage value.

Depletion base = $640000 - $0 = $640000

Estimated recoverable units = 500000 tons

Putting the values in the unit depletion rate formula, we get,

Units depletion Rate = $640000 / 500000 = $1.28 per ton

In the next step, we will compute the depreciation expense as per below:

Depreciation expense = Unit depletion rate * No. of units extracted

No. of units extracted was 80000 tons, so

Depletion = $1.28 * 80000 = $102400

Required journal entry for depletion expense of 2019 is:

Debit Depletion expense $102400

Credit Accumulated depletion     $102400